Please help me solve the equation 6+2(3j+-2)=4(1+j)
Answer by nyc_function(2741)

6+ 2(3j +(-2))= 4( 1 + j)
6 + 6j - 4 = 4 + 4j
6j + 2 = 4 + 4j
2 - 4 = 4j - 6j
-2 = -2j
-2/-2 = j
2/2 = j
1 = j
Done!